OpenStudy (anonymous):

Write an equation of a line that passes through the point (-5, 7) and whose slope is 6.?

OpenStudy (anonymous):

y = mx + b, where m is your slope and b is your y intercept. The y intercept fot this is 37, so the equation is y = 6x + 37

OpenStudy (anonymous):

so would it be y+5=6[x+7] ?

OpenStudy (anonymous):

Do you need it in point slope form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Because if you do, it's y + 5 = 6(x + 7) :) you are right :)

OpenStudy (anonymous):

lol thank you